OpenStudy (jacobbirddog):

Help fan+medallion! Which is the contrapositive of the conditional statement? If an animal is a deer, then it has four legs. A. If an animal has four legs, then it is a deer. B. If an animal does not have four legs, then it has four legs. C. If an animal does not have four legs, then it is not a deer. D. If an animal is not a deer, then it is does not have four legs.

OpenStudy (anonymous):

wait no

OpenStudy (jacobbirddog):

yep. i have one more and this last one is screwed up

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

um i would go with A. because all the others are crazy!!

OpenStudy (jacobbirddog):

Which shows the contrapositive of the conditional statement, and if the conditional statement and the contrapositive are true or false? If x – 7 ≠ 10, then x ≠ 17. A. If x – 7 = 10, then x = 17. The conditional statement and the contrapositive are both true. B. If x ≠ 17, then x – 7 ≠ 10. The conditional statement and the contrapositive are both false. C. If x = 17, then x – 7 = 10. The conditional statement is false and the contrapositive is true. D. If x = 17, then x – 7 = 10. The conditional statement and the contrapositive are both true.

OpenStudy (anonymous):

Sorry man i have no clue that is way to confusing.

OpenStudy (jacobbirddog):

ikr O_O

OpenStudy (anonymous):

Have a great day man see you later :)

OpenStudy (jacobbirddog):

U2